NEAR-INFRARED OVERTONE SPECTROSCOPY OF TRITIATED WATER

\underline{KAORI KOBAYASHI},TOMOYA ENOKIDA, DAISUKE IIO, YUTA YAMADA, Department of Physics, University of Toyama, 3190 Gofuku, Toyama, 930-8555 Japan; MASANORI HARA, YUJI HATANO, Hydrogen Isotope Research Center, University of Toyama, 3190 Gofuku, Toyama, 930-8555 Japan.

The tritiated water molecules (here HTO and T_{2}O) are important isotopomers of water. The spectroscopic information is essential from the viewpoint of the basic science. However, the high-resolution spectroscopic studies of tritiated water were limited probably due to the radioactive nature of tritium. The microwave studies of these species were carried out and the molecular constants of the ground state were determined. ^{a} ^{,}^{b} The fundamental ν_{1} band and the ν_{3} band of HTO were reported. ^{c} ^{,}^{d} For T_{2}O, the ν_{3} band was studied. ^{e} At 1.3 micron region, overtone and combination bands are expected. In this study we prepared tritiated water of high concentration which is necessary for the near-infrared measurement and carried out frequency modulated near-infrared spectroscopy. Many lines which are not listed in HITRAN were observed. We will report the current status of the analysis.

aF. C. De Lucia, P. Helminger, W. Gordy, H. W. Morgan and P. A. Staats, Phys. Rev. A \underline{8}, 2785 (1973).

bP. Helminger, F. C. De Lucia, W. Gordy, P. A. Staats and H. W. Morgan, Phys. Rev. A \underline{10}, 1072 (1974).

cS. D. Cope, D. K. Russell, H. A. Fry, L. H. Jones, J. E. Barefield, J. Mol. Spectrosc. \underline{127}, 464 (1988).

dM. Tine and L. H. Coudert, International Symposium on Molecular Spectroscopy, 66th Meeting, RE11 (2005).

eS. D. Cope, D. K. Russell, H. A. Fry, L. H. Jones, J. E. Barefield, J. Mol. Spectrosc. \underline{120}, 311 (1986).